OpenStudy (anonymous):

Find the domain f(x)=(x+2)/(x^2-1)

OpenStudy (astrophysics):

You're looking at the restrictions, so where the denominator is = 0, so solve for \[x^2 - 1 = 0\]

OpenStudy (anonymous):

\[\pm1\]

OpenStudy (anonymous):

i got that part. i don't exactly know how to write it in interval notation tho

OpenStudy (anonymous):

(\[-\infty,-1]\cup(0,1]\cup(1,\infty)\]

OpenStudy (anonymous):

i don't think i did it right

OpenStudy (astrophysics):

You could write it as such \[( - \infty, - 1) \cup (1, \infty)\] because open bracket implies it's not included, a square bracket means, it would be included in the domain.

OpenStudy (astrophysics):

of course I forgot 0 xD

OpenStudy (astrophysics):

But you get the point right?

OpenStudy (usukidoll):

I used to hate those write in set notation or interval notation parts to the question because I got confused with ( and ]. This was before OpenStudy popped up.

OpenStudy (astrophysics):

Yeah I can't remember last time I even used interval notation to be honest

OpenStudy (usukidoll):

College Algebra had truckloads... but the thing is it doesn't appear again afterwards.

OpenStudy (astrophysics):

Yea, I used it in calculus 1, then never saw it again haha...

OpenStudy (astrophysics):

 so this is how it looks like in a number line (random example) \[(- \infty, -1/2] \cup [3, \infty)\] I think this will make it clear on what exactly is going on.

OpenStudy (astrophysics):